Ref. 1046146:
0.95 = 0.4v + 0.4
0.95 - 0.4 = 0.4v + 0.4 - 0.4
0.55 = 0.4v
0.55 / 0.4 = 0.4v / 0.4
1.375 = v
Where am I correct ?
Also, I have read that if, an amount of water, in this case, the amt. of alcohol is diluted, but when alcohol is added, the amt. remains unchanged. Am I on the right track ?



Answer by Alan3354(69443)

0.55 = 0.4v
0.55 / 0.4 = 0.4v / 0.4
1.375 = v
Where am I correct ?
Also, I have read that if, an amount of water, in this case, the amt. of alcohol is diluted, but when alcohol is added, the amt. remains unchanged. Am I on the right track ?
===========================================
A 1 qt. mixture of alcohol and water is 95% alcohol. The mixture is to be reduced to 40% alcohol. Determine amount of distilled water to be added.
--------------
Add W quarts of water
---
1*95 + W*0 = (1 + W)* 40
40W + 40 = 95
40W = 55
W = 11/8 quarts